# The ghost-gluon vertex in the presence of the Gribov horizon: general kinematics

### v1

@@ III.2 The soft gluon limit
The calculation of the ghost-gluon vertex for any momentum requires the calculation of the diagrams shown in fig. 1. Since this calculation is rather long to be performed manually, in a previous work some of us considered the soft-gluon limit of $\Gamma_{A\bar{c}c}$, i.e., the limit in which the gluon momentum $k\rightarrow 0$ Mintz et al. (2018).
In the soft gluon limit, each diagram takes a simplified form and furthermore, since diagram (IV) of fig. 1 is proportional to the gluon momentum (and it is finite), it vanishes for $k\rightarrow 0$. In this limit, the one-loop correction for the vertex, which has been explicitly calculated in Mintz et al. (2018), reads
@@
which are the same in the RGZ framework as in perturbative YM theory Taylor (1971). These are direct consequences of the Ward identities of the action (II).
@@ IV.1 Feynman diagrams
$D_{(I)}=-ig^{3}\frac{N}{2}f^{abc}p_{\nu}(p+k)_{\rho}\int\frac{d^{d}q}{(2\pi)^{d}}q_{\mu}P^{\perp}_{\nu\rho}(q+p)D_{AA}(q+p)D_{\bar{c}c}(q-k)D_{\bar{c}c}(q)\,,$
$D_{(II)}=-ig^{3}\frac{N}{2}f^{abc}p_{\eta}(p+k)_{\omega}\int\frac{d^{d}q}{(2\pi)^{d}}\left(-2k_{\rho}\delta_{\mu\nu}+2k_{\nu}\delta_{\mu\rho}+(2q-k)_{\mu}\delta_{\nu\rho}\right)P^{\perp}_{\nu\eta}(q)P^{\perp}_{\rho\omega}(q-k)D_{AA}(q)D_{AA}(q-k)D_{\bar{c}c}(q+p)\,,$
$D_{(III)}=ig^{3}\frac{N^{2}}{4}f^{abc}(g\gamma^{2})^{2}p_{\nu}(p+k)_{\rho}\int\frac{d^{d}q}{(2\pi)^{d}}(2q-k)_{\mu}P^{\perp}_{\rho\eta}(q-k)P^{\perp}_{\nu\eta}(q)D_{A\varphi}(q-k)D_{A\varphi}(q)D_{\bar{c}c}(q+p)\,,$
$f^{ade}D_{(IV.1)}=ig^{3}\frac{N^{2}}{4}f^{abc}(g\gamma^{2})p_{\eta}k_{\omega}(p+k)_{\sigma}\int\frac{d^{d}q}{(2\pi)^{d}}P^{\perp}_{\sigma\mu}(q-k)P^{\perp}_{\eta\omega}(q)D_{A\varphi}(q-k)D_{AA}(q)D_{\bar{c}c}(q+p)\,,$
$f^{ade}D_{(IV.2)}=-ig^{3}\frac{N^{2}}{4}f^{abc}(g\gamma^{2})p_{\sigma}k_{\omega}(p+k)_{\eta}\int\frac{d^{d}q}{(2\pi)^{d}}P^{\perp}_{\sigma\mu}(q)P^{\perp}_{\eta\omega}(q-k)D_{A\varphi}(q)D_{AA}(q-k)D_{\bar{c}c}(q+p)\,,$ (29)
@@
The above Feynman integrals depend on two external momenta, p and k, and feature only one Lorentz index which is not contracted, $\mu$. As a result, all of them can be expressed as $F(p^{2},k^{2},p\cdot k,M^{2},m^{2},\lambda)p_{\mu}+H(p^{2},k^{2},p\cdot k,M^{2},m^{2},\lambda)k_{\mu}$, where $F$ and $H$ are scalar integrals. In the first stage of our computation we found the corresponding scalar integrals for each diagram [^3] The contributions for the dressing function $G^{CCA}$ come from the $p_{\mu}$ component entirely. and evaluated the color factors by means of the ColorMath package Sjödahl (2013).
In a second stage, we reduced the scalar integrals to one-loop master integrals Passarino and Veltman (1979) of the type
@@
We implemented the reduction to master integrals using an algorithm in Mathematica based on the Fire package Smirnov and Chuharev (2020). The reduction for each diagram is presented in a supplementary material. The analysis is performed for an arbitrary number of colors $N$ with the purpose of comparing with alternative approaches, both for the SU(2) and SU(3) gauge groups.
Once the reduction is finished we need to compute the one-loop master integrals. For both $A$ and $B$ there are well-known analytic results (see, for instance, ref. Martin (2003)). In the case of the master integral $C$, when needed, we evaluated it numerically.
@@ IV.2 Kinematic configurations
In this paper we aim at extending the analysis of the ghost-gluon vertex within the RGZ framework, initially explored in Mintz et al. (2018), beyond the soft gluon limit, to include arbitrary kinematic configurations. We focus on the specific kinematic setups for which we have access to lattice simulations, namely the symmetric and orthogonal configurations.
@@ IV.3 Numerical values of the RGZ parameters
In principle, the three massive parameters could be determined self-consistently using the Gribov gap equation and minimizing the effective potential of the theory with respect to two condensates. Here, however, we shall fix these parameters using lattice input. The parameters $M,m$ and $\lambda$ at tree-level were determined for SU(2) in Cucchieri et al. (2012) and for SU(3) in Oliveira and Silva (2012), by fitting the lattice data of the gluon propagator from YM theory with an analytic expression of the form:
@@
SU(2) - Ref.Cucchieri et al. (2012) 2.51 -1.92 1.52
SU(3) - Ref.Oliveira and Silva (2012) 4.47 -3.77 2.04
Table 1: RGZ parameters fitted from lattice results for SU(2) Cucchieri et al. (2012) and SU(3) Oliveira and Silva (2012).
@@ IV.4 Toy model of the running coupling
Firstly, we simply investigate at a qualitative level the variation of the dressing function $G^{CCA}$ as we change fixed values of $g$. As we will see, this is enough to reproduce either the IR or the UV but not both regimes at the same time. Most likely this is due to the presence of large logarithms in the UV domain, which spoils the validity of perturbation theory. This problem can be overcome by introducing the running of the various parameters of the RGZ framework and choosing a renormalization scale of the same order of the external momentum: $\mu\sim p$. To achieve this we would need to evaluate the two-point functions of the theory by introducing an infrared-safe scheme, of the type introduced in Pelaez et al. (2013) for instance.
@@ V.1 SU(2)
We begin by analyzing the ghost-gluon vertex dressing function $G^{CCA}$ for the SU(2) gauge group in the symmetric and orthogonal kinematic configurations, as well as in the case of the soft gluon limit. The latter case was previously investigated in Mintz et al. (2018). However, for completeness and because the analysis we perform in this article differs slightly from the one developed in Mintz et al. (2018), we include those results here as well.
@@ V.1.1 Fitting the ghost-gluon vertex dressing function
Figure 2: The SU(2) scalar function $G^{CCA}$ as a function of the antighost external momentum in three distinct configurations: the symmetric (top left), orthogonal (top right) and the soft gluon limit (bottom) for various values of a fixed gauge coupling. Blue and black points refer to lattice data extracted from refs. Cucchieri et al. (2008) and Maas (2020), respectively.
@@
Figure 4: The scalar function $G^{CCA}$ for the SU(2) case as a function of the antighost external momentum in three distinct configurations: the symmetric (top left), orthogonal (top right) and the soft gluon limit (bottom). The red curve represents the RGZ fit of the lattice data by using the toy model of eq. 38 for the gauge coupling. The values of the parameters, $\Lambda=2.75$ GeV and $g_{0}=3.9$, are chosen to minimize the joint relative error $\chi_{\text{joint}}$. Regarding the symmetric and orthogonal configurations, the blue curve is the prediction from the CF model at one-loop order Pelaez et al. (2013). In the soft gluon limit, the CF results are global fits of the gluon and ghost propagators along with the dressing function $G^{CCA}$ at one- and two-loop order Barrios et al. (2020). Blue and black points refer to lattice data extracted from refs. Cucchieri et al. (2008) and Maas (2020), respectively.
@@ V.1.2 Contribution of each diagram to the $G^{CCA}$ fit
We note that genuinely RGZ diagrams, (III) and (IV), represent a small contribution at one-loop order as compared to the largest contribution of diagram (I) and, to a lesser degree, diagram (II). This is in agreement with the results obtained by CF model Tissier and Wschebor (2010, 2011); Pelaez et al. (2013); Peláez et al. (2014, 2015); Gracey et al. (2019); Peláez et al. (2021); Figueroa and Peláez (2022); Barrios et al. (2020, 2021, 2022) where only QCD-like diagrams with massive gluons are taken into account. Interestingly, the effect of diagrams (III) and (IV) is a reduction of the peak of $G^{CCA}$, at around $p=1.5$ GeV.
@@ V.2.1 Fitting the ghost-gluon vertex dressing function
In fig. 6 we see $G^{CCA}$ as a function of the antighost momentum for the symmetric and orthogonal configurations as well as in the soft gluon limit. Similarly to what we observed in the case of the SU(2) gauge group, as for the soft gluon limit it seems that we are capable of reproducing the lattice data by using a fixed value of the gauge coupling $g$, as was already pointed out in Mintz et al. (2018).
Figure 6: The scalar function $G^{CCA}$ for the SU(3) case as a function of the antighost external momentum in three distinct configurations: the symmetric (top left), orthogonal (top right) and the soft gluon limit (bottom) for various values of a fixed gauge coupling. The lattice data were extracted from the plots in Ilgenfritz et al. (2007) using WebPlotDigitizer Rohatg . We estimated the error of the extraction procedure to be at most 0.8%.
Figure 6: The scalar function $G^{CCA}$ for the SU(3) case as a function of the antighost external momentum in three distinct configurations: the symmetric (top left), orthogonal (top right) and the soft gluon limit (bottom) for various values of a fixed gauge coupling. The lattice data were extracted from the plots in Ilgenfritz et al. (2007) using WebPlotDigitizer Rohatg . We estimated the error of the extraction procedure to be at most 0.8%.
@@
Figure 8: The scalar function $G^{CCA}$ for the SU(3) case as a function of the antighost external momentum in three distinct configurations: the symmetric (top left), orthogonal (top right) and the soft gluon limit (bottom). The red curve corresponds to the RGZ fit of the lattice data, by utilizing the toy model of eq. 38 for the gauge coupling. Parameters $\Lambda=0.70$ GeV and $g_{0}=2.15$ were chosen to minimize the joint relative error with the lattice data in the case of the soft gluon limit. Lattice data are extracted from Ilgenfritz et al. (2007). We compare the results with other approaches in the continuum. Specifically with $G^{CCA}$ obtained from DSE, considering the bare (DSE - I) and the dressed three-gluon vertex (DSE - II) Aguilar et al. (2019) and $G^{CCA}$ from the CF model at one- and two-loop order Pelaez et al. (2013); Barrios et al. (2020).
We note that our results for $G^{CCA}$ are consistent with the outcomes of the CF model Pelaez et al. (2013); Barrios et al. (2020) and DSE Aguilar et al. (2019). Furthermore, the RGZ fit shows a strong agreement with simulations, resembling the results obtained with a fixed value of the gauge coupling. Interestingly, while the UV tails of the DSE and CF curves overlap in the UV, the UV tails from the RGZ curves fall more rapidly in comparison.
@@ V.2.2 Contribution of each diagram to the $G^{CCA}$ fit
We observe that in the case of the symmetric and orthogonal configurations the behaviour of $G^{CCA}$ is dominated by diagram (II), while the contribution of diagram (I) is counterbalanced by the contributions of the genuine RGZ diagrams (III) and (IV). Overall, the relative weights of the these contributions do not significantly differ from the SU(2) case.
Figure 9: Contribution of each diagram to $G^{CCA}$ in three distinct configurations: the symmetric (top left), orthogonal (top right) and the soft gluon limit (bottom), using the gauge coupling described by eq. 38. Parameters, $\Lambda=0.70$ GeV and $g_{0}=2.15$, are chosen to minimize the relative error $\chi$ in the soft gluon limit. The convention of colors and line-styles in the orthogonal and soft configurations is the same as the symmetric case. The solid, red curve is the sum of all diagrams. The numbering of the diagrams is provided in fig. 1. Diagram (IV) refers to the full contribution $\frac{2g\gamma^{2}f^{ade}}{k^{2}+M^{2}}\Gamma_{\bar{c}^{b}c^{c}\varphi^{de}_{\mu}}(k,p,r)$.

### v2

@@ III.2 The soft gluon limit
The calculation of the ghost-gluon vertex for any momentum requires the calculation of the diagrams shown in fig. 1. Since this calculation is rather long to be performed manually, in a previous work some of us considered the soft-gluon limit of $\Gamma_{A\bar{c}c}$, i.e., the limit in which the gluon momentum $k\rightarrow 0$ [1].
In the soft gluon limit, each diagram takes a simplified form. Moreover, diagrams (IV.1) and (IV.2) vanish. In reality, as for the soft gluon limit, these contributions vanish at all orders of perturbation theory. In the case of $\Gamma_{\bar{c}^{b}c^{c}\varphi_{\mu}^{de}}$, this is due to the derivative nature of the $\varphi\bar{\varphi}A$ vertex and the transversality of the gluon propagator, as well as all propagators involving the $\varphi$ field (see appendix A). This property could also be inferred by looking at the function $\Gamma_{\bar{c}^{b}c^{c}\bar{\varphi}_{\mu}^{de}}$ which is identical to $\Gamma_{\bar{c}^{b}c^{c}\varphi_{\mu}^{de}}$. As $\Gamma_{\bar{c}^{b}c^{c}\bar{\varphi}_{\mu}^{de}}$ is always proportional to the momentum of the external gluon, it vanishes as this momentum goes to zero [^3] In both cases of $\Gamma_{\bar{c}^{b}c^{c}\varphi_{\mu}^{de}}$ and $\Gamma_{\bar{c}^{b}c^{c}\bar{\varphi}_{\mu}^{de}}$ we are considering loop corrections to be regular as $k\to 0$. This is indeed the case of the RGZ framework, owing to the fact that the parameters $m$, $M$ and $\gamma$ act as natural IR regulators of the theory..
In this limit, the one-loop correction for the vertex, which has been explicitly calculated in [1], reads
@@
which are the same in the RGZ framework as in perturbative YM theory [64]. These are direct consequences of the Ward identities of the action (II).
@@ IV.1 Feynman diagrams
$D_{\text{(I)}}=-ig^{3}\frac{N}{2}f^{abc}p_{\nu}(p+k)_{\rho}\int\frac{d^{d}q}{(2\pi)^{d}}q_{\mu}P^{\perp}_{\nu\rho}(q+p)D_{AA}(q+p)D_{\bar{c}c}(q-k)D_{\bar{c}c}(q)\,,$
$D_{\text{(II)}}=-ig^{3}\frac{N}{2}f^{abc}p_{\eta}(p+k)_{\omega}\int\frac{d^{d}q}{(2\pi)^{d}}\left(-2k_{\rho}\delta_{\mu\nu}+2k_{\nu}\delta_{\mu\rho}+(2q-k)_{\mu}\delta_{\nu\rho}\right)P^{\perp}_{\nu\eta}(q)P^{\perp}_{\rho\omega}(q-k)D_{AA}(q)D_{AA}(q-k)D_{\bar{c}c}(q+p)\,,$
$D_{\text{(III)}}=ig^{3}\frac{N^{2}}{4}f^{abc}(g\gamma^{2})^{2}p_{\nu}(p+k)_{\rho}\int\frac{d^{d}q}{(2\pi)^{d}}(2q-k)_{\mu}P^{\perp}_{\rho\eta}(q-k)P^{\perp}_{\nu\eta}(q)D_{A\varphi}(q-k)D_{A\varphi}(q)D_{\bar{c}c}(q+p)\,,$
$f^{ade}D_{\text{(IV.1)}}=ig^{3}\frac{N^{2}}{4}f^{abc}(g\gamma^{2})p_{\eta}k_{\omega}(p+k)_{\sigma}\int\frac{d^{d}q}{(2\pi)^{d}}P^{\perp}_{\sigma\mu}(q-k)P^{\perp}_{\eta\omega}(q)D_{A\varphi}(q-k)D_{AA}(q)D_{\bar{c}c}(q+p)\,,$
$f^{ade}D_{\text{(IV.2)}}=-ig^{3}\frac{N^{2}}{4}f^{abc}(g\gamma^{2})p_{\sigma}k_{\omega}(p+k)_{\eta}\int\frac{d^{d}q}{(2\pi)^{d}}P^{\perp}_{\sigma\mu}(q)P^{\perp}_{\eta\omega}(q-k)D_{A\varphi}(q)D_{AA}(q-k)D_{\bar{c}c}(q+p)\,,$ (29)
@@
The above Feynman integrals depend on two external momenta, p and k, and feature only one Lorentz index which is not contracted, $\mu$. As a result, all of them can be expressed as $F(p^{2},k^{2},p\cdot k,M^{2},m^{2},\lambda)p_{\mu}+H(p^{2},k^{2},p\cdot k,M^{2},m^{2},\lambda)k_{\mu}$, where $F$ and $H$ are scalar integrals. In the first stage of our computation we found the corresponding scalar integrals for each diagram [^5] The contributions for the dressing function $G^{CCA}$ come from the $p_{\mu}$ component entirely. and evaluated the color factors by means of the ColorMath package [65].
In a second stage, we reduced the scalar integrals to one-loop master integrals [66] of the type
@@
We implemented the reduction to master integrals using an algorithm in Mathematica based on the Fire package [67]. The reduction for each diagram is presented in a supplementary material. The analysis is performed for an arbitrary number of colors $N$ with the purpose of comparing with alternative approaches, both for the SU(2) and SU(3) gauge groups.
Once the reduction is finished we need to compute the one-loop master integrals. For both $A$ and $B$ there are well-known analytic results (see, for instance, ref. [68]). In the case of the master integral $C$, when needed, we evaluated it numerically.
@@ IV.2 Kinematic configurations
In this paper we aim at extending the analysis of the ghost-gluon vertex within the RGZ framework, initially explored in [1], beyond the soft gluon limit, to include arbitrary kinematic configurations. We focus on the specific kinematic setups for which we have access to lattice simulations, namely the symmetric and orthogonal configurations.
@@ IV.3 Numerical values of the RGZ parameters
In principle, the three massive parameters could be determined self-consistently using the Gribov gap equation and minimizing the effective potential of the theory with respect to two condensates. Here, however, we shall fix these parameters using lattice input. The parameters $M,m$ and $\lambda$ at tree-level were determined for SU(2) in [69] and for SU(3) in [70], by fitting the lattice data of the gluon propagator from YM theory with an analytic expression of the form:
@@
SU(2) - Ref.[69] 2.51 -1.92 1.52
SU(3) - Ref.[70] 4.47 -3.77 2.04
Table 1: RGZ parameters fitted from lattice results for SU(2) [69] and SU(3) [70].
@@ IV.4 Toy model of the running coupling
Firstly, we simply investigate at a qualitative level the variation of the dressing function $G^{CCA}$ as we change fixed values of $g$. As we will see, this is enough to reproduce either the IR or the UV but not both regimes at the same time. Most likely this is due to the presence of large logarithms in the UV domain, which spoils the validity of perturbation theory. This problem can be overcome by introducing the running of the various parameters of the RGZ framework and choosing a renormalization scale of the same order of the external momentum: $\mu\sim p$. To achieve this we would need to evaluate the two-point functions of the theory by introducing an infrared-safe scheme, of the type introduced in [29] for instance.
@@ V.1 SU(2)
We begin by analyzing the ghost-gluon vertex dressing function $G^{CCA}$ for the SU(2) gauge group in the symmetric and orthogonal kinematic configurations, as well as in the case of the soft gluon limit. The latter case was previously investigated in [1]. However, for completeness and because the analysis we perform in this article differs slightly from the one developed in [1], we include those results here as well.
@@ V.1.1 Fitting the ghost-gluon vertex dressing function
Figure 2: The SU(2) scalar function $G^{CCA}$ as a function of the antighost external momentum in three distinct configurations: the symmetric (top left), orthogonal (top right) and the soft gluon limit (bottom) for various values of a fixed gauge coupling. Blue and black points refer to lattice data extracted from refs. [71] and [72], respectively.
@@
Figure 4: The scalar function $G^{CCA}$ for the SU(2) case as a function of the antighost external momentum in three distinct configurations: the symmetric (top left), orthogonal (top right) and the soft gluon limit (bottom). The red curve represents the RGZ fit of the lattice data by using the toy model of eq. 38 for the gauge coupling. The values of the parameters, $\Lambda=2.75$ GeV and $g_{0}=3.9$, are chosen to minimize the joint relative error $\chi_{\text{joint}}$. Regarding the symmetric and orthogonal configurations, the blue curve is the prediction from the CF model at one-loop order [29]. In the soft gluon limit, the CF results are global fits of the gluon and ghost propagators along with the dressing function $G^{CCA}$ at one- and two-loop order [35]. Blue and black points refer to lattice data extracted from refs. [71] and [72], respectively.
@@ V.1.2 Contribution of each diagram to the $G^{CCA}$ fit
We note that genuinely RGZ diagrams, (III) and (IV), represent a small contribution at one-loop order as compared to the largest contribution of diagram (II) and, to a lesser degree, diagram (I). This is in agreement with the results obtained by CF model [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37] where only QCD-like diagrams with massive gluons are taken into account. Interestingly, as for the symmetric and orthogonal configurations, the effect of diagrams (III) and (IV) is a reduction of the peak of $G^{CCA}$, at around $p=1.5$ GeV. In the case of the soft gluon limit, we observe that the contributions coming from diagrams (I) and (II) are of similar order whereas the contribution of diagram (IV) vanishes, as already pointed out in section III.2.
@@ V.2.1 Fitting the ghost-gluon vertex dressing function
In fig. 6 we see $G^{CCA}$ as a function of the antighost momentum for the symmetric and orthogonal configurations as well as in the soft gluon limit. Similarly to what we observed in the case of the SU(2) gauge group, as for the soft gluon limit it seems that we are capable of reproducing the lattice data by using a fixed value of the gauge coupling $g$, as was already pointed out in [1].
Figure 6: The scalar function $G^{CCA}$ for the SU(3) case as a function of the antighost external momentum in three distinct configurations: the symmetric (top left), orthogonal (top right) and the soft gluon limit (bottom) for various values of a fixed gauge coupling. The lattice data were extracted from the plots in [73] using WebPlotDigitizer [74]. We estimated the error of the extraction procedure to be at most 0.8%.
@@
Figure 8: The scalar function $G^{CCA}$ for the SU(3) case as a function of the antighost external momentum in three distinct configurations: the symmetric (top left), orthogonal (top right) and the soft gluon limit (bottom). The red curve corresponds to the RGZ fit of the lattice data, by utilizing the toy model of eq. 38 for the gauge coupling. Parameters $\Lambda=0.70$ GeV and $g_{0}=2.15$ were chosen to minimize the joint relative error with the lattice data in the case of the soft gluon limit. Lattice data are extracted from [73]. We compare the results with other approaches in the continuum. Specifically with $G^{CCA}$ obtained from DSE, considering the bare (DSE - I) and the dressed three-gluon vertex (DSE - II) [15] and $G^{CCA}$ from the CF model at one- and two-loop order [29, 35].
We note that our results for $G^{CCA}$ are consistent with the outcomes of the CF model [29, 35] and DSE [15]. Furthermore, the RGZ fit shows a strong agreement with simulations, resembling the results obtained with a fixed value of the gauge coupling. Interestingly, while the UV tails of the DSE and CF curves overlap in the UV, the UV tails from the RGZ curves fall more rapidly in comparison.
@@ V.2.2 Contribution of each diagram to the $G^{CCA}$ fit
We observe that in the case of the symmetric and orthogonal configurations the behaviour of $G^{CCA}$ is dominated by diagram (II), while the contribution of diagram (I) is counterbalanced by the contributions of the genuine RGZ diagrams (III) and (IV). In the case of the soft gluon limit, diagrams (I) and (II) both significantly contribute to the function $G^{CCA}$, whereas diagram (IV) vanishes, as explained in section III.2. Overall, the relative weights of these contributions do not significantly differ from the SU(2) case.
Figure 9: Contribution of each diagram to $G^{CCA}$ for the SU(3) case in three distinct configurations: the symmetric (top left), orthogonal (top right) and the soft gluon limit (bottom), using the gauge coupling described by eq. 38. Parameters, $\Lambda=0.70$ GeV and $g_{0}=2.15$, are chosen to minimize the relative error $\chi$ in the soft gluon limit. The convention of colors and line-styles in the orthogonal and soft configurations is the same as the symmetric case. The solid, red curve is the sum of all diagrams. The numbering of the diagrams is provided in fig. 1. Diagram (IV) refers to the full contribution $\frac{2g\gamma^{2}f^{ade}}{k^{2}+M^{2}}\Gamma_{\bar{c}^{b}c^{c}\varphi^{de}_{\mu}}(k,p,r)$.